OpenStudy (anonymous):

Plz simplify the fraction :- 22386 -------- 9900

OpenStudy (yttrium):

maybe try to reduce both the numerator and the denominator into their prime factors.

OpenStudy (anonymous):

i wanna have the answer cuz i want to check if the answer given in the book is right or wrong

OpenStudy (yttrium):

whatwhat's the answer on the book?

OpenStudy (wolf1728):

Reducing both numbers to their prime factors: 22,386 = 2*3*7*13*41 9,900 = 2*2*3*3*5*5*11 So both numbers only have 2 and 3 in common so we can divide both numbers by 6 and we get: -7*13*41 divided by 2*3*5*5*11 equals -3,731 divided by 1,650

OpenStudy (anonymous):

or \[-2\frac{ 431 }{ 1650 }\]